# Hamiltonian Learning using Machine Learning Models Trained with Continuous Measurements

### v1

@@ Abstract
We build upon recent work on using Machine Learning models to estimate Hamiltonian parameters using continuous weak measurement of qubits as input. We consider two settings for the training of our model: (1) supervised learning where the weak measurement training record can be labeled with known Hamiltonian parameters, and (2) unsupervised learning where no labels are available. The first has the advantage of not requiring an explicit representation of the quantum state, thus potentially scaling very favorably to larger number of qubits. The second requires the implementation of a physical model to map the Hamiltonian parameters to a measurement record, which we implement using an integrator of the physical model with a recurrent neural network to provide a model-free correction at every time step to account for small effects not captured by the physical model. We test our construction on a system of two qubits and demonstrate accurate prediction of multiple physical parameters in both the supervised and unsupervised context. We demonstrate that the model benefits from larger training sets establishing that it is in fact “learning,” and we show robustness to errors in the assumed physical model by achieving accurate parameter estimation in the presence of unanticipated single particle relaxation.
preprint: SAND2024-04128O
@@ I Introduction
In this work, we build on recent advances in device characterization applying ML models to the continuous measurement of qubits [20, 21, 19]. In contrast with the parameter estimation approach of Ref. [19], where a stochastic master equation is used as a trainable model, our ML model learns a direct map from continuous measurement inputs to the system parameters of interest. In the case where the measurement records are associated with known device parameters (the case of supervised learning), this approach has two advantages. First, the model is completely independent of any state representation. The measurements need only provide enough information to estimate the parameters of interest and not the complete quantum state, potentially freeing it from the curse of dimensionality that comes from the exponential scaling of the dimension of the Hilbert space and thus making it more scalable.
@@ II Physical System
where $\{X_{i},Y_{i},Z_{i}\}$ is the set of Pauli operators for qubit $i\in\left\{1,2\right\}$, $\mathcal{D}[L](\rho)=L\rho L^{\dagger}-\frac{1}{2}\left\{L^{\dagger}L\rho+\rho L^{\dagger}L\right\}$ is the Lindblad super-operator, $\mathcal{H}[L](\rho)=L\rho+\rho L^{\dagger}-\rho\mbox{Tr}\left[\rho\left(L+L^{\dagger}\right)\right]$ is the measurement super-operator, $\eta=0.1469$ is the efficiency of the measurement, and $L_{i}=\sqrt{\kappa}C_{i},$ where $C_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is the weak measurement operator with $\kappa=3.326$ radians/$\mu$s. Here we have suppressed the dependence on $t$ for $\rho$, and we have adopted the normalization convention $\hbar=1$. The parameter values are selected to be comparable to those used in Ref. [19].
@@
Figure 1: The physical system of two qubits in a microwave cavity subject to a Rabi drive $\Omega$, single particle relaxation with rate $\gamma_{s}$, and dephasing due to weak measurement back-action with rate $\kappa$
@@ III.2 Model Description
Our objective is to train a ML model to estimate the physical master equation parameters in Eq. (II) by observing the weak measurement records described by Eq. (2). Input data is provided for training of the ML model in the form of a collection of averaged weak measurement records, where the average is over some number of records $d$ to reduce the effects of diffusive noise. The output of the model is either (a) the physical master equation parameters of interest in the supervised case where labeled training data is available, or (b) the predicted measurement record satisfying the unconditioned master equation (Eq. (II)) in the unsupervised case where true parameters are not available. The former configuration can be interpreted as an encoder mapping noisy trajectories to estimated parameter values, while the latter configuration adds a decoder that takes the estimated parameters as input and produces a predicted solution to the unconditioned master equation. The full architecture is shown in Fig. 3 and can be viewed as a denoising autoencoder taking noisy input trajectories and producing output with the noise removed.
@@
where $\theta$ is a set of ODE parameters, $f_{\text{model}}$ is the drift term of the form of Eq. (II), and $f_{\text{LSTM}}$ is an LSTM cell and as such is a non-linear, model free function of the time, state, and parameters.
@@
It should be noted that while the integrator used in the decoder for this work is an Euler integrator for the unconditioned master equation, this piece is completely modular in design and can be replaced with any integrator for simulating quantum dynamics. For example, a TDVP integrator using MPS could be used for larger systems that are approximated well by a MPS ansatz, with the trainable decoder parameters accounting for errors introduced by this approach.
@@ III.3 Model Interpretation
Figure 3: Diagram of the full ML model. Voltage records are measured from the cavity system producing noisy averaged voltage records (i), these are sent through a neural network encoder (ii) producing system parameters as output (iii). These parameters are used by the flex integrator decoder (iv) to produce noise-free voltage estimates (v)
@@ III.4 Training and Data
In order to generate the data for training, validation, and testing, we use an Euler-Maruyama integrator to solve Eq. (2) for both qubits for each of $40$ values of $\epsilon$ evenly spaced on $[0,2)$ with fixed $\Omega=1.395$ radians/$\mu$s, and the same number of $\Omega$ values evenly spaced on $[1,5)$ with fixed $\epsilon=1.0$ radians/$\mu$s for a total of $K=80$ ($\Omega$, $\epsilon$) pairs. For each pair of values ($\Omega_{k}$, $\epsilon_{k}$), $N=32{,}000$ measurement trajectories are simulated for $T=4\mu s$. Half of the collection of pair of values ($\Omega_{k}$, $\epsilon_{k}$), corresponding to every other parameter pair, is used for training. The other half, corresponding to parameters midway between training values, is split evenly to be used for validation and testing. This ensures the training set contains a disjoint set of parameter pairs from the validation/test set. The two endpoints of the validation/test set are excluded from each end to ensure that the training data extends slightly beyond the domain of validation/test values. Only weak measurement records are being used, and no strong measurements are being performed.
@@
where $p$ is the dimension of $\theta_{k}$, the number of parameters being estimated. In the unsupervised case, the loss function is the MSE between each $x_{k}$ and the estimated clean measurement record
@@
During evaluation, validation and test error is evaluated in a similar manner to how training mini-batches are selected, with groups of size $d$ randomly chosen from the full validation and test sets, followed by the calculation of the MSE. This process is repeated 100 times and the average is taken as the calculated MSE for each set.
@@ IV.1 Supervised Learning
Figure 5: Estimations of the parameter pair $\theta_{k}=(\Omega_{k},\epsilon_{k})$ (top) and the squared error (bottom) averaged over four noisy trajectories for a model trained on clean data and noisy data, then evaluated on noisy data. Shuffle evaluated test set MSEs are (8.67e-3, 5.83e-4) and (4.94e-3, 4.87e-4) respectively
@@ IV.3 Model Correction
In this section we demonstrate the ability of the decoder to correct for dynamics not explicitly considered in the physical model (Eq. (1a)). This is done by enabling training for the decoder LSTM parameters. For the datasets, we simulate 30,000 trajectories with fixed $\Omega=1.395$ radians/$\mu$s but with varying $\epsilon$, which is the parameter to be estimated, while adding a dissipative term for each qubit corresponding to the Lindblad operator $\mathcal{D}[\sqrt{\gamma_{s}}\sigma_{i}^{-}](\rho)$ in Eq. (1a), where $\gamma_{s}=0.1$ and $\sigma_{i}^{-}=\frac{1}{2}(X_{i}-iY_{i})$ is the single particle relaxation operator mapping the excited state to the ground state.
@@
Figure 6: Estimate of $\epsilon$ and $P(Y_{i}=0)$ when $\epsilon=1.7$ and measuring $X$ with spin-up in $Z$ as the initial state for both qubits for training data simulated with single particle relaxation rate $\gamma_{s}=0.1$ but a model that does not explicitly account for $\gamma_{s}$. Results are shown for the physical model alone and with a correction learned by the decoder LSTM
@@
5,000 2.84e-3 0.142 3.82e-3
$\infty$ 2.12e-6 0.145 2.71e-4
Table 5: MSE of $\epsilon$ estimates using different $\gamma_{s}$ values in the model. The final column shows results when learning is enabled for the free parameters in the decoder to account for the unanticipated term $\gamma_{s}$. Measurement time spacing is $\Delta t=2^{-8}\mu s$
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation the Quantum Leap Big Idea under Grant No. OMA-1936388. AKR and CM acknowledge support from NSF QLCI OMA-2016244. This article has been co-authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan.

### v2

@@ Abstract
We build upon recent work on using Machine Learning models to estimate Hamiltonian parameters using continuous weak measurement of qubits as input. We consider two settings for the training of our model: (1) supervised learning where the weak measurement training record can be labeled with known Hamiltonian parameters, and (2) unsupervised learning where no labels are available. The first has the advantage of not requiring an explicit representation of the quantum state, thus potentially scaling very favorably to a larger number of qubits. The second requires the implementation of a physical model to map the Hamiltonian parameters to a measurement record, which we implement using an integrator of the physical model with a recurrent neural network to provide a model-free correction at every time step to account for small effects not captured by the physical model. We test our construction on a system of two qubits and demonstrate accurate prediction of multiple physical parameters in both the supervised and unsupervised contexts. We demonstrate that the model benefits from larger training sets establishing that it is in fact “learning,” and we show robustness to errors in the assumed physical model by achieving accurate parameter estimation in the presence of unanticipated single particle relaxation.
@@ I Introduction
In this work, we build on recent advances in device characterization applying ML models to the continuous measurement of qubits [20, 21, 19]. In contrast with the parameter estimation approach of Ref. [19], where a stochastic master equation is used as a trainable model, our ML model learns a direct map from continuous measurement inputs to the system parameters of interest. In the case where the measurement records are associated with known device parameters (the case of supervised learning), this approach has two advantages. First, the model is completely independent of any state representation (in contrast with approaches like Ref. [22]). The measurements need only provide enough information to estimate the parameters of interest and not the complete quantum state, potentially freeing it from the curse of dimensionality that comes from the exponential scaling of the dimension of the Hilbert space and thus making it more scalable.
@@
Finally, we contrast our work with recent methods for Hamiltonian learning based on Gibbs state measurements or real-time evolution [22, 31, 32, 33, 34, 35, 36, 37, 38]. In our work the model is trained and performs predictions based on continuous measurements modeled by a stochastic master equation rather than Hamiltonian or deterministic master equation evolution combined with strong measurements. Our approach that is applicable to both Hamiltonian and Lindblad parameters requires that the parameters to be learned have an observable impact on the continuous measurements, and it does not use the preparation of a steady state or energy eigenstate [22, 32, 38] or changing the dynamics by adaptively changing the Hamiltonian [36] or introducing additional dynamics [34, 35]. It is likely adaptive approaches could enhance our learning as well. Finally, the discrepancy modeling in our approach is unique in that it allows for the combination of knowledge about the physical system with model-free machine learning to enhance estimation accuracy and ML model interpretability.
@@ II Physical System
where $\{X_{i},Y_{i},Z_{i}\}$ is the set of Pauli operators for qubit $i\in\left\{1,2\right\}$, $\mathcal{D}[L](\rho)=L\rho L^{\dagger}-\frac{1}{2}\left\{L^{\dagger}L\rho+\rho L^{\dagger}L\right\}$ is the Lindblad super-operator, $\mathcal{H}[L](\rho)=L\rho+\rho L^{\dagger}-\rho\mbox{Tr}\left[\rho\left(L+L^{\dagger}\right)\right]$ is the measurement super-operator, $\eta$ is the efficiency of the measurement, and $L_{i}=\sqrt{\kappa}C_{i},$ where $C_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is the weak measurement operator. Here we have suppressed the dependence on $t$ for $\rho$, and we have adopted the normalization convention $\hbar=1$.
@@
Figure 1: The physical system of two qubits in a microwave cavity subject to a Rabi drive $\Omega$, single particle relaxation with rate $\gamma_{s}$, and dephasing due to weak measurement back-action with rate $\kappa$.
@@ III.2 Model Description
Our objective is to train a ML model to estimate the physical master equation parameters in Eq. (1) by observing the weak measurement records described by Eq. (2). Input data is provided for training of the ML model in the form of a collection of averaged weak measurement records, where the average is over some number of records $d$ to reduce the effects of diffusive noise. The output of the model is either (a) the physical master equation parameters of interest in the supervised case where labeled training data is available, or (b) the predicted measurement record satisfying the unconditioned master equation (Eq. (3)) in the unsupervised case where true parameters are not available. The former configuration can be interpreted as an encoder mapping noisy trajectories to estimated parameter values, while the latter configuration adds a decoder that takes the estimated parameters as input and produces a predicted solution to the unconditioned master equation. The full architecture is shown in Fig. 3 and can be viewed as a denoising autoencoder taking noisy input trajectories and producing output with the noise removed.
@@
where $\theta$ is a set of ODE parameters, $f_{\text{model}}$ is the drift term of the form of Eq. (3), and $f_{\text{LSTM}}$ is an LSTM cell and as such is a non-linear, model free function of the time, state, and parameters. While we are assuming for this work, as discussed in Sec. III.4, that the parameters of the measurement equation in Eq. (3b) are known with high precision, this approach could easily be extended to this equation as well.
@@
It should be noted that while the integrator used in the decoder for this work is an Euler integrator for the unconditioned master equation, this piece is completely modular in design and can be replaced with any integrator for simulating quantum dynamics. For example, a TDVP integrator using MPS could be used for larger systems that are approximated well by a MPS ansatz, with the trainable decoder parameters accounting for errors introduced by this approach. It should also be noted that, since the output of the decoder is the solution to the unconditioned master equation, it contains an ODE rather than an SDE integrator. Once trained, however, it would be possible to switch over to an SDE integrator, for example by switching from Euler to Euler-Maruyama, by simulating noise. The model would then serve as a generative model that could be useful for simulation purposes.
@@ III.3 Model Interpretation
Figure 3: Diagram of the full ML model. Voltage records are measured from the cavity system producing noisy averaged voltage records (i), these are sent through a neural network encoder (ii) producing system parameters as output (iii). These parameters are used by the flex integrator decoder (iv) to produce noise-free voltage estimates (v).
@@ III.4 Training and Data
In order to generate the data for training, validation, and testing, we use an Euler-Maruyama integrator to solve Eq. (2) for both qubits for each of $40$ values of $\epsilon$ evenly spaced on $[0,2)$ with fixed $\Omega=1.395$ radians/$\mu$s, and the same number of $\Omega$ values evenly spaced on $[1,5)$ with fixed $\epsilon=1.0$ radians/$\mu$s for a total of $K=80$ ($\Omega$, $\epsilon$) pairs. For each pair of values ($\Omega_{k}$, $\epsilon_{k}$), $N=32{,}000$ measurement trajectories are simulated for $T=4\mu s$. Half of the collection of pair of values ($\Omega_{k}$, $\epsilon_{k}$), corresponding to every other parameter pair, is used for training. The other half, corresponding to parameters midway between training values, is split evenly in the number of trajectories to be used for validation and testing. This ensures the training set contains a disjoint set of parameter pairs from the validation/test set. The two endpoints of the validation/test set are excluded from each end to ensure that the training data extends slightly beyond the domain of validation/test values. Only weak measurement records are being used, and no strong measurements are being performed.
Values of $\kappa=3.326$ radians/$\mu$s and $\eta=0.1469$ are selected to be consistent with the example of superconducting qubits as found, for example, in Ref. [19], though with a stronger measurement back-action $\kappa$ after simulations showed a larger value to offer a good balance between dephasing rate and measurement noise. While either $\Omega$, $\epsilon$, or both are treated as unknown parameters and comprise the latent space of our model, it is assumed in our simulations that $\kappa$ and $\eta$ are known with high precision as they have already been calibrated. Our approach can be used to perform this calibration step, since the latent space of the model is not limited to Hamiltonian parameters and could support any combination of Hamiltonian and Lindblad parameters as unknown variables. Alternatively, other calibration methods found in the literature can be used to estimate them, such as in [52]. It should be noted, however, that in practice if calibration errors are present in the model values, the degrees of freedom in the decoder should make parameter estimation robust to these errors as demonstrated in what is likely the more challenging case of unanticipated single particle relaxation examined in Sec. IV.3. The impact on parameter estimation accuracy is expected to be small for small calibration errors. A study quantifying this dependency is an interesting area for future work.
@@
where $p$ is the dimension of $\theta_{k}$, the number of parameters being estimated. This is further averaged over the $M$ groups of size $d$ when multiple trajectory groups are provided as input. In the unsupervised case, the loss function is the MSE between each $x_{k}$ and the estimated clean measurement record
@@
During evaluation, validation and test error is evaluated in a similar manner to how training mini-batches are selected, with groups of size $d$ randomly chosen from the full validation and test sets, followed by the calculation of the MSE. This process is repeated 100 times and the average is taken as the calculated MSE for each set. Note that for some scenarios only one group will be available as input, either because the data is noise free or because the group size equals the dataset size, in which case the shuffling has no effect.
@@ IV.1 Supervised Learning
Figure 5: Estimations of the parameter pair $\theta_{k}=(\Omega_{k},\epsilon_{k})$ (top) and the squared error (bottom) averaged over four noisy trajectory groups of $d=4{,}000$ measurement records for a model trained on clean data and noisy data, then evaluated on noisy data. Shuffle evaluated test set MSEs are (8.67e-3, 5.83e-4) and (4.94e-3, 4.87e-4) respectively.
@@ IV.3 Model Correction
In this section we demonstrate the ability of the decoder to correct for dynamics not explicitly considered in the physical model (Eq. (1a)). This is done by enabling training for the decoder LSTM parameters. For the datasets, we simulate 30,000 trajectories with fixed $\Omega=1.395$ radians/$\mu$s that is known to the model, but with varying unknown $\epsilon$ which is the parameter to be estimated. We add a dissipative term for each qubit corresponding to the Lindblad operator $\mathcal{D}[\sqrt{\gamma_{s}}\sigma_{i}^{-}](\rho)$ in Eq. (1a), where $\gamma_{s}=0.1$ and $\sigma_{i}^{-}=\frac{1}{2}(X_{i}-iY_{i})$ is the single particle relaxation operator mapping the excited state to the ground state. Note that while the choice of a single unknown parameter is made in this section to provide a clear and straightforward demonstration, the model still supports multiple possible parameters to be estimated in addition to the decoder degrees of freedom.
@@
Figure 6: Estimate of $\epsilon$ (top) and $P(Y_{i}=0)$ when $\epsilon=1.7$ (bottom) and measuring $X$ with spin-up in $Z$ as the initial state for both qubits for training data simulated with single particle relaxation rate $\gamma_{s}=0.1$ but a model that does not explicitly account for $\gamma_{s}$. Results are shown for the physical model alone and with a correction learned by the decoder LSTM.
@@
5,000 2.84e-3 0.142 3.59e-3
$\infty$ 2.12e-6 0.145 3.26e-4
Table 5: MSE of $\epsilon$ estimates using different $\gamma_{s}$ values in the model. The final column shows results when learning is enabled for the free parameters in the decoder to account for the unanticipated term $\gamma_{s}$. Measurement time spacing is $\Delta t=2^{-8}\mu s$.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation the Quantum Leap Big Idea under Grant No. OMA-1936388. AKR and CM acknowledge support from NSF QLCI OMA-2016244. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. This article has been co-authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan.